C6H6(l) + Br2(l) ----> C6H5Br(l) + HBr(g)


If 60.5 g of benzene is mixed with 134g of bromine,


How many grams of bromobenzene are formed in the reaction?


Mass=...........g



Please write it with the steps.

Thanks.....


1
Expert's answer
2021-02-26T06:02:13-0500

From the chemical equation;

"60.5gC_6H_5\times" "1mol C_6H_6\over 78.11gC_6H_6" "\times" "1mol C_6H_5Br\over 1mol C_6H_6" "\times" "157.0gC_6H_5Br\over 1mol C_6H_5Br" "=121.60gC_6H_5Br"

"134gBr_2\times" "1mol Br_2\over 159.6gBr_2" "\times" "1mol C_6H_5Br\over 1mol Br_2" "\times" "157.0gC_6H_5Br\over 1mol C_6H_5Br" "=131.81gC_6H_5Br"


Mass"=253.41g"